# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ VI-B.7 BD-RIS with Sensing and ISAC
The future networks are expected to support an increasing number of spectrum-demanding applications, which will cause increasing radio spectrum congestion. This growth motivates the design of shared paradigms, which enable cooperative spectrum and resource sharing among systems. One such solution is the emerging ISAC technique [6], which realizes dual functions with shared hardware, platform, and resources to reduce costs and increase resource utilization. However, since wireless sensing relies on strong LoS channels, which can be obstructed by obstacles in complex wireless propagation environments, it might be difficult to achieve a satisfactory ISAC performance in practical scenarios. To tackle this challenge, [141] has proposed a multi-sector BD-RIS self-sensing system, where an active source controller is installed on one sector such that signals can be scattered toward other sectors to achieve full-space coverage. BD-RIS can also perform as a bridge to link active and passive localization [142] and effectively increase localization precision.
Beyond enabling sensing and localization, there have been a few works on BD-RIS-aided ISAC [143, 144, 145, 146, 147, 148]. Specifically, [143] proposed a hybrid BD-RIS-aided ISAC system where targets and users are located at both transmitting and reflecting sectors of BD-RIS, such that the performance of both functions and coverage is improved. In addition, deploying BD-RIS in ISAC systems can also help to increase signal-to-noise ratio (SNR) for both functions [144], reduce transmit power [145], enhance network throughput [146], achieve better communication-sensing trade-off [147], and meet ISAC outage requirements [148].
VI-B.8 BD-RIS for WPT and SWIPT
@@
Energy harvesting techniques are fundamental enablers for realizing seamless and green connectivity between low-power devices and supporting IoT in future wireless networks [160]. Among various energy harvesting techniques, WPT is a paradigm making full use of wireless to deliver energy [7]. Beyond WPT, a shared paradigm has been proposed, namely SWIPT [7, 28], which enables using wireless to not only deliver energy but also convey information to best use the RF spectrum. To further enlarge the effective distance of WPT/SWIPT and improve the power transfer efficiency, [149, 150] have deployed BD-RIS in the wireless system, showing that BD-RIS can flexibly shape the wireless channel to facilitate WPT. In addition, [151] has focused on a SWIPT system and demonstrated that BD-RIS is beneficial in improving both harvested energy of energy users and spectral efficiency of information users.
VI-B.9 BD-RIS Aiding Mobile Edge Computing (MEC)
@@
Mobile edge computing (MEC) [161] is a platform that pushes mobile computing, network control, and storage to the network edges, such as base stations and access points. In this way, MEC enables resource-limited mobile devices to complete computational-intensive tasks with much reduced latency. Due to these benefits, MEC supports broad mobile applications, such as augmented reality services, IoT applications, and connected vehicles. To further improve the scalability and performance of MEC to meet the growing user demands in 5G and future 6G networks, [152] has proposed deploying BD-RIS in the wireless propagation environment as an enabler to significantly improve the number of completed task bits.
VII Challenges and Future Research Directions of BD-RIS
@@ VII-A BD-RIS Implementation
Figure 35: Photograph of the prototype of the hybrid BD-RIS with $4\times 4$ cells and the controlling systems [162].
The hardware implementation of BD-RIS is a fundamental yet challenging issue. Until now, there has been the implementation of one special type of BD-RIS, namely STAR-RIS (or IOS, STARS) [37]. In addition, a hybrid mode BD-RIS has been recently implemented using power splitters as shown in Fig. 35, enabling independent beam control of reflected and transmitted waves [162]. Furthermore, a reflective mode BD-RIS, specifically a tridiagonal BD-RIS, has been implemented through a tunable load network allowing interconnections among adjacent RIS elements as shown in Fig. 36 [163]. These results show that it is physically feasible to implement BD-RIS using existing RF techniques, designs, and devices. However, the comprehensive implementation of various BD-RIS modes and architectures is still on its way. Based on the illustration in Section III, an $M$-element BD-RIS consists of an $M$-antenna array and an $M$-port reconfigurable impedance network.
  • On the one hand, the implementation of the $M$-antenna array varies with operation modes. For the reflective mode, the conventional uniform linear or planar array can be directly used. For the hybrid mode, each pair of antennas with uni-directional radiation pattern (such as patch antennas) should be back to back placed to construct a cell, and all cells should be arranged in a uniform array. For the multi-sector mode, every $L$ antennas with corresponding beamwidth should be placed as a polygon and all cells should be arranged in a uniform array.
  • On the other hand, the implementation of the $M$-port reconfigurable impedance network varies with architectures. For fully/group-, tree/forest-, stem/band-connected architectures, as per the circuit topology, varactors consisting of inductance and tunable capacitance can be used to implement continuous-value reconfigurable admittance components; PIN diodes can be used to implement discrete-value reconfigurable admittance components. For dynamically connected architectures, additional switches are required between every two elements. For non-reciprocal architectures, non-reciprocal devices, such as circulators, isolators, and gyators, are required and the number of them depends on specific circuit designs.
@@
Figure 36: Photograph and schematic of the prototype of the tridiagonal BD-RIS with $8$ elements [163].
VII-A.1 Challenge
@@
The unique challenge of implementing BD-RIS compared to D-RIS lies in the increasing circuit complexity, cost, and power losses of the reconfigurable impedance network. Especially when the number of elements and/or groups increases, the required number of reconfigurable admittance components significantly increases. This will cause a heavy burden on circuit design and, more importantly, accumulated power losses induced by more reconfigurable admittance components, which may weaken the performance benefits of BD-RIS architectures.
VII-A.2 Future Research Direction
@@
Although the tree/forest-connected architecture [40] and the stem/band-connected architecture [41] have been proposed to achieve the best performance-complexity trade-off in, respectively, MISO and multi-user MIMO systems, these conclusions still stop in the analytical layer. The practical limitation of these architectures taking into account the circuit design feasibility, cost, and power losses still remains unexplored. Therefore, it is challenging but worthwhile to implement and prototype different BD-RIS architectures to have a practical picture on which kind of architecture could best balance the performance and hardware considerations.
VII-B Active BD-RIS
@@
The existing literature of BD-RIS has focused purely on its passive (at best lossless) form, in the sense that the power scattered by the surface is no larger than the incident power. This assumption comes from D-RIS literature and brings many advantages to both D-RIS and BD-RIS, such as low cost, low power consumption, and negligible thermal noise [8]. However, since the array gain provided by numerous elements of RIS cannot fully compensate for the huge multiplicative fading induced by RIS, the performance improvement of passive RIS is very limited. For D-RIS, the gain in some scenarios is even not visible when the direct link between the transmitter and receiver exists. For BD-RIS, in SISO systems, the fully-connected architecture achieves a maximum of 62% gain over D-RIS [26] with the absence of the direct link, which is not sufficiently high, especially when the direct link is strong. To break the fundamental performance bottleneck caused by multiplicative fading, the concept of active D-RIS has been proposed in [21]. The key feature of such D-RIS is its ability to reflect signals with amplified powers, thanks to the introduction of power amplifiers in RIS elements. The benefits of active D-RIS has also been shown in many aspects, such as achieving a substantial sum-rate gain of 130% [21], achieving higher energy efficiency than passive D-RIS [164], and further enhancing both communication quality and sensing performance [165]. This motivates the consideration of active BD-RIS: Will integrating active power amplifiers in BD-RIS architectures provide orders of gains?
VII-B.1 Challenge
@@
To answer the above question, one fundamental yet challenging issue is to establish a physics-consistent model for active BD-RIS. Different from passive BD-RIS which is modeled as an antenna array connected to a multi-port reconfigurable impedance network, the active BD-RIS, following the modeling principle in [21], should be modeled as a series connection of an antenna array, a multi-port reconfigurable impedance network, and a reflection power amplifier network. In this sense, the admittance/impedance/scattering matrices themselves of the reconfigurable impedance network are not sufficient to characterize the active BD-RIS, and the impact of power amplifiers should be carefully taken into account.
VII-B.2 Future Research Direction
@@
The consideration of making BD-RIS active opens many interesting and meaningful research directions, such as theoretically deriving the power scaling law of active BD-RIS for different channel fading conditions; revisiting the optimal performance-circuit complexity Pareto frontier or exploring the optimal performance-power consumption Pareto frontier for active BD-RIS; and fully making use of the active property to reduce channel estimation error and overhead. Beyond these examples, other fundamental directions in passive BD-RIS, such as the impact of hardware impairments, will also be worth studying in active BD-RIS. In this case, hardware impairments will not be limited to those summarized in Section V, but will also include the imperfections of power amplifiers.
VII-C Artificial Intelligence (AI)-Driven Beamforming Solutions
@@
From the beamforming design perspective, most existing literature on BD-RIS uses conventional optimization methods to design BD-RIS beamforming, aiming at either finding closed-form solutions with a guaranteed performance or finding iterative solutions with a guaranteed convergence. However, these optimization methods have the following drawbacks: 1) They depend heavily on accurate physics modeling of wireless channels and thus are not easy to be used in practical scenarios where theoretical models cannot fully reflect the true characteristics of wireless environments; 2) The computational complexity can be extremely high when considering large-dimensional scenarios, which are unfortunately quite common in the modern world. Although BD-RIS theoretically has appealing benefits supported by these optimization methods, it is important to improve the scalability and feasibility of BD-RIS design methods to prepare for the possible future commercialization of BD-RIS. This thus motivates the consideration of using artificial intelligence (AI)-driven solutions, since they are not sensitive to models and suitable for large-scale problems [166, 167].
VII-C.1 Challenge
@@
VII-C.2 Future Research Direction
Although [68, 69] have adopted deep reinforcement learning (DRL) and meta-learning to design BD-RIS-aided wireless communication systems and shown the benefits of such AI-driven solutions in being friendly to large-scale wireless systems, the study of AI-enabled BD-RIS-aided wireless systems is still in its infancy. Future research avenues include, but are not limited to, designing more efficient online adaptation schemes to enable highly reliable instantaneous transmissions and establishing proper AI models that can properly adapt to various BD-RIS architectures, that are robust to real hardware constraints of BD-RIS, and that have good training overhead-performance trade-offs.
VII-D CSI-Free Protocols
@@ VII-D.1 Challenge
In D-RIS-aided systems, protocols based on beam training schemes have been proposed to reduce overhead and avoid the requirement for accurate CSI acquisition [169, 170, 171]. Specifically, a codebook consisting of possible scattering matrices of RIS is predefined in an offline stage. During the online stage, the transmitter consecutively sends pilot signals while the RIS switches between codewords in the codebook. The receiver collects the received signal powers corresponding to all codewords, finds the index of the codeword which leads to the best received power, and feeds this index back to the transmitter. In this way, the channel estimation and beamforming phases in the typically used protocol are replaced with a beam training phase supported by a predefined codebook, such that the RIS can be optimized without explicitly knowing the CSI. The main difficulty in this protocol lies in the construction of RIS codebook. Especially when it comes to BD-RIS with various architectures, it is challenging to construct codebooks whose codewords perfectly follow the constraints of corresponding architectures and have sufficiently large diversity to reflect the flexibility of BD-RIS architectures compared to D-RIS.
VII-D.2 Future Research Direction

### v2

@@ VII-B BD-RIS for Sensing and ISAC
Future networks are expected to support an increasing number of spectrum-demanding applications, which will cause increasing radio spectrum congestion. This growth motivates the design of shared paradigms, which enable cooperative spectrum and resource sharing among systems. One such solution is the emerging ISAC technique [liu2022integrated], which realizes dual functions with shared hardware, platform, and resources to reduce costs and increase resource utilization. However, since wireless sensing relies on strong LoS channels, which can be obstructed by obstacles in complex wireless propagation environments, it might be difficult to achieve a satisfactory ISAC performance in practical scenarios. To tackle this challenge, [zhang2024full] has proposed a multi-sector BD-RIS self-sensing system, where an active source controller is installed on one sector such that signals can be scattered toward other sectors to achieve full-space coverage. BD-RIS can also perform as a bridge to link active and passive localization [raeisi2024efficient] and effectively increase localization precision. Beyond enabling sensing and localization, there have been a few works on BD-RIS-aided ISAC [wang2024dual, esmaeilbeig2024beyond, guang2024power, liu2024enhancing, chen2024transmitter, nguyen2025beyond]. Specifically, [wang2024dual] proposed a hybrid BD-RIS-aided ISAC system where targets and users are located at both transmitting and reflecting sectors of BD-RIS, such that the performance of both functions and coverage is improved. In addition, deploying BD-RIS in ISAC systems can also help to increase signal-to-noise ratio (SNR) for both functions [esmaeilbeig2024beyond], reduce transmit power [guang2024power], enhance network throughput [liu2024enhancing], achieve better communication-sensing trade-off [chen2024transmitter], and meet ISAC outage requirements [nguyen2025beyond].
VII-C BD-RIS for WPT and SWIPT
@@
Energy harvesting techniques are fundamental enablers for realizing seamless and green connectivity between low-power devices and supporting IoT in future wireless networks [lopez2021massive]. Among various energy harvesting techniques, WPT is a paradigm making full use of wireless to deliver energy [clerckx2021wireless]. Beyond WPT, a shared paradigm has been proposed, namely SWIPT [clerckx2021wireless, perera2017simultaneous], which enables using wireless to not only deliver energy but also convey information to best use the RF spectrum. To further enlarge the effective distance of WPT/SWIPT and improve the power transfer efficiency, [azarbahram2025beyond, azarbahram2025beamforming] have deployed BD-RIS in the wireless system, showing that BD-RIS can flexibly shape the wireless channel to facilitate WPT. In addition, [hua2024cell] has focused on a SWIPT system and demonstrated that BD-RIS is beneficial in improving both harvested energy of energy users and spectral efficiency of information users.
VII-D BD-RIS with Other Techniques and Systems
VII-D.1 BD-RIS and Stacked Intelligent Metasurface (SIM)
Stacked intelligent surface (SIM) [an2023stacked] consists in stacking multiple layers of RIS or metasurfaces to provide more flexible wave manipulations. While SIM layers have been implemented using D-RIS [an2023stacked], [nerini2024physically] has recently shown that SIM layers can be implemented using BD-RIS. Theoretical and numerical results have demonstrated that 1-layer SIM implemented with BD-RIS (tree-connected) is sufficient to achieve performance upper-bound in SISO systems, while multi-layer SIM implemented with D-RIS is suboptimal.
VII-D.2 BD-RIS with Advanced Multiple Access
Multiple access refers to techniques that make use of resource dimensions, such as time, frequency, space, and power, to serve multiple users. Non-orthogonal multiple access (NOMA) has been actively studied recently due to its benefits in exploiting the available resources more efficiently with the aid of superposition coding at the transmitter and successive interference cancellation (SIC) at the receiver [liu2017nonorthogonal]. To further boost system performance, the interaction between NOMA and BD-RIS has been studied in [zhang2024beyond, agarwal2025fairness], showing that better user fairness can be achieved. In addition, rate-splitting multiple access (RSMA) is a novel and general framework for non-orthogonal transmissions, softly including NOMA as an extreme case [mao2022rate] and moving beyond that to provide enhanced performance. Given that channel acquisition of RIS-aided wireless systems is generally challenging while RSMA is robust to CSI imperfections, the integration of the two has emerged [fang2022fully, kim2023group, soleymani2023optimization, khisa2024gradient, li2024synergizing]. Specifically, [fang2022fully, kim2023group, soleymani2023optimization, khisa2024gradient] have focused mainly on perfect CSI and have shown the benefits of such an integration in increasing spectral efficiency, fairness rate, and fairness energy efficiency. [li2024synergizing] has further shown that the integration of the two can still increase system sum-rate when only imperfect CSI is available at the transmitter.
VII-D.3 BD-RIS Aiding Unmanned Aerial Vehicle (UAV) and Non-Terrestrial Networks (NTN)
Unmanned Aerial Vehicle (UAV) has gained attention for 5G and 6G networks due to its advantages in extending coverage, easing the deployment, and supporting controllable mobility [gupta2015survey]. These advantages of UAVs can be further strengthened by cooperating with BD-RIS. [huroon2023optimized] and [lin2024securing] studied the transmission strategy and user scheduling schemes of a ground-based BD-RIS in a network consisting of multiple UAVs, showing that the system performance can be much improved with the aid of BD-RIS. [mahmood2023joint] proposed to colocate BD-RIS and UAVs to minimize the computation latency and UAV hovering time. Not limited to UAV-enabled systems, BD-RIS can also provide performance benefits in general non-terrestrial networks (NTN) typically realized by satellites, high-altitude platform stations (HAPS), and UAVs [azari2022evolution]. Results in [khan2025enhancing, khan2025beyondnew] have shown that having a BD-RIS mounted UAV [khan2025enhancing] or HAPS [khan2025beyondnew] as a secondary transmitter in a cognitive radio enabled NTN helps increase secrecy rate and spectral efficiency.
VII-D.4 BD-RIS Aiding Mobile Edge Computing (MEC)
@@
Mobile edge computing (MEC) [chen2022irs] is a platform that pushes mobile computing, network control, and storage to the network edges, such as base stations and access points. In this way, MEC enables resource-limited mobile devices to complete computational-intensive tasks with much reduced latency. To further improve the scalability and performance of MEC to meet the growing user demands in 5G and future 6G networks, [qin2025joint] has proposed deploying BD-RIS in the wireless propagation environment as an enabler to significantly improve the number of completed task bits.
VIII Challenges and Future Research Directions of BD-RIS
@@ VIII-A BD-RIS Implementation
Figure 35: Photograph of the prototype of the hybrid BD-RIS with $4\times 4$ cells and the controlling systems [ming2025hybrid].
@@
Figure 36: Photograph and schematic of the prototype of the tridiagonal BD-RIS with 88 elements [tap2025].
The hardware implementation of BD-RIS is a fundamental yet challenging issue. Until now, there has been the implementation of one special type of BD-RIS, namely STAR-RIS (or IOS, STARS) [zhang2022intelligent]. In addition, a hybrid mode BD-RIS has been recently implemented using power splitters as shown in Fig. 35, enabling independent beam control of reflected and transmitted waves [ming2025hybrid]. Furthermore, a reflective mode BD-RIS, specifically a tridiagonal BD-RIS, has been implemented through a tunable load network allowing interconnections among adjacent RIS elements as shown in Fig. 36 [tap2025]. These results show that it is physically feasible to implement BD-RIS using existing RF techniques, designs, and devices. However, the comprehensive implementation of various BD-RIS modes and architectures is still on its way. Based on the illustration in Section III, an $M$-element BD-RIS consists of an $M$-antenna array and an $M$-port reconfigurable impedance network.
  • The implementation of the $M$-antenna array varies with operation modes. For the reflective mode, the conventional uniform linear or planar array can be directly used. For the hybrid mode, each pair of antennas with uni-directional radiation pattern (such as patch antennas) should be back to back placed to construct a cell, and all cells should be arranged in a uniform array. For the multi-sector mode, every $L$ antennas with corresponding beamwidth should be placed as a polygon and all cells should be arranged in a uniform array.
  • The implementation of the $M$-port reconfigurable impedance network varies with architectures. For fully/group-, tree/forest-, stem/band-connected architectures, as per the circuit topology, varactors consisting of inductance and tunable capacitance can be used to implement continuous-value reconfigurable admittance components; PIN diodes can be used to implement discrete-value reconfigurable admittance components. For dynamically connected architectures, additional switches are required between every two elements. For non-reciprocal architectures, non-reciprocal devices, such as circulators, isolators, and gyators, are required and the number of them depends on specific circuit designs.
VIII-A.1 Challenge
@@
The unique challenge of implementing BD-RIS compared to D-RIS lies in the increasing circuit complexity, cost, and power losses of the reconfigurable impedance network. Especially when the number of elements and/or groups increases, the required number of reconfigurable admittance components significantly increases. This will cause a heavy burden on circuit design and, more importantly, accumulated power losses, and power consumption of drive circuits and elements induced by more reconfigurable admittance components [wang2024power], which may weaken the performance benefits of BD-RIS architectures. In addition, for future system-level experiments, the theoretical channel estimation and beamforming optimization methods might require too much time such that an effective data transmission cannot be guaranteed.
VIII-A.2 Future Research Direction
@@
Although the tree/forest-connected architecture [nerini2024beyond] and the stem/band-connected architecture [wu2025beyond] have been proposed to achieve the best performance-complexity trade-off in, respectively, MISO and multi-user MIMO systems, these conclusions still stop in the analytical layer. The practical limitation of these architectures taking into account the circuit design feasibility, cost, and power losses still remains unexplored. In addition, it is important to establish a fair power consumption model for BD-RIS architectures and validate its feasibility by practical measurements. How to balance the circuit complexity, computational complexity for optimization, and system performance for a practical BD-RIS-assisted test bed is also a crucial problem. Therefore, it is challenging but worthwhile to implement and prototype different BD-RIS architectures to have a practical picture on which kind of architecture could best balance the performance, hardware considerations, and computational complexity.
VIII-B Active BD-RIS
@@
The existing literature of BD-RIS has focused purely on its passive (at best lossless) form, in the sense that the power scattered by the surface is no larger than the incident power. This assumption comes from D-RIS literature and brings many advantages to both D-RIS and BD-RIS, such as low cost, low power consumption, and negligible thermal noise [di2020smart]. However, since the array gain provided by numerous elements of RIS cannot fully compensate for the huge multiplicative fading induced by RIS, the performance improvement of passive RIS is very limited. For D-RIS, the gain in some scenarios is even not visible when the direct link between the transmitter and receiver exists. For BD-RIS, in SISO systems, the fully-connected architecture achieves a maximum of 62% gain over D-RIS [shen2021] with the absence of the direct link, which is not sufficiently high, especially when the direct link is strong. To break the fundamental performance bottleneck caused by multiplicative fading, the concept of active D-RIS has been proposed in [zhang2022active]. The key feature of such D-RIS is its ability to reflect signals with amplified powers, thanks to the introduction of power amplifiers in RIS elements. The benefits of active D-RIS has also been shown in many aspects, such as achieving a substantial sum-rate gain of 130% [zhang2022active], achieving higher energy efficiency than passive D-RIS [zhi2022active], and further enhancing both communication quality and sensing performance [zhu2023joint]. This motivates the consideration of active BD-RIS: Will integrating active power amplifiers in BD-RIS architectures provide orders of gains?
VIII-B.1 Challenge
@@
To answer the above question, one fundamental yet challenging issue is to establish a physics-consistent model for active BD-RIS. Different from passive BD-RIS which is modeled as an antenna array connected to a multi-port reconfigurable impedance network, the active BD-RIS, following the modeling principle in [zhang2022active], should be modeled as a series connection of an antenna array, a multi-port reconfigurable impedance network, and a reflection power amplifier network. In this sense, the admittance/impedance/scattering matrices themselves of the reconfigurable impedance network are not sufficient to characterize the active BD-RIS, and the impact of power amplifiers should be carefully taken into account.
VIII-B.2 Future Research Direction
@@
The consideration of making BD-RIS active opens many interesting and meaningful research directions, such as theoretically deriving the power scaling law of active BD-RIS for different channel fading conditions; revisiting the optimal performance-circuit complexity Pareto frontier or exploring the optimal performance-power consumption Pareto frontier for active BD-RIS; and fully making use of the active property to reduce channel estimation error and overhead. Beyond these examples, other fundamental directions in passive BD-RIS, such as the impact of hardware impairments, will also be worth studying in active BD-RIS. In this case, hardware impairments will not be limited to those summarized in Section VI, but will also include the imperfections of power amplifiers.
VIII-C Artificial Intelligence (AI)-Driven Beamforming Solutions
@@
From the beamforming design perspective, most existing literature on BD-RIS uses conventional optimization methods to design BD-RIS beamforming, aiming at either finding closed-form solutions with a guaranteed performance or finding iterative solutions with a guaranteed convergence. However, these optimization methods have the following drawbacks: 1) They depend heavily on accurate physics modeling of wireless channels and thus are not easy to be used in practical scenarios where theoretical models cannot fully reflect the true characteristics of wireless environments; 2) The computational complexity can be extremely high when considering large-dimensional scenarios, which are unfortunately quite common in the modern world. Although BD-RIS theoretically has appealing benefits supported by these optimization methods, it is important to improve the scalability and feasibility of BD-RIS design methods to prepare for the possible future commercialization of BD-RIS. This thus motivates the consideration of using artificial intelligence (AI)-driven solutions, since they are not sensitive to models and suitable for large-scale problems [eldar2022machine, jiang2016machine].
VIII-C.1 Challenge
@@
VIII-C.2 Future Research Direction
Although [sobhi2024joint, loli2024meta] have adopted deep reinforcement learning (DRL) and meta-learning to design BD-RIS-aided wireless communication systems and shown the benefits of such AI-driven solutions in being friendly to large-scale wireless systems, the study in this direction is still in its infancy. Future research avenues include, but are not limited to, designing more efficient online adaptation schemes to enable highly reliable instantaneous transmissions and establishing proper AI models that can properly adapt to various BD-RIS architectures, that are robust to real hardware constraints of BD-RIS, and that have good training overhead-performance trade-offs.
VIII-D CSI-Free Protocols
@@ VIII-D.1 Challenge
In D-RIS-aided systems, protocols based on beam training schemes have been proposed to reduce overhead and avoid the requirement for accurate CSI acquisition [mu2024reconfigurable, wang2022beam, liu2023low]. Specifically, a codebook consisting of possible scattering matrices of RIS is predefined in an offline stage. During the online stage, the transmitter consecutively sends pilot signals while the RIS switches between codewords in the codebook. The receiver collects the received signal powers corresponding to all codewords, finds the index of the codeword which leads to the best received power, and feeds this index back to the transmitter. In this way, the channel estimation and beamforming phases in the typically used protocol are replaced with a beam training phase supported by a predefined codebook, such that the RIS can be optimized without explicitly knowing the CSI. The main difficulty in this protocol lies in the construction of RIS codebook. Especially when it comes to BD-RIS with various architectures, it is challenging to construct codebooks whose codewords perfectly follow the constraints of corresponding architectures and have sufficiently large diversity to reflect the flexibility of BD-RIS architectures compared to D-RIS.
VIII-D.2 Future Research Direction